Suppose K= 9.4 x 10^-3 at a certain temperature for the reaction PCl5--->PCl3+Cl2. Iif it is found that the concentration of PCl5 is twice the concentration of PCl3, what must be the concentration of Cl2 under these conditions?
Answer #1

Hi,

The equilibrium constant fro this reaction would be

K = [PCl3] [Cl2] / [PCl5]

9.4 * 10^-3 = x [Cl2] / 2x

[Cl2] = 9.4 * 10^-3 * 2 = 18.8 * 10^-3